OpenStudy (anonymous):

im willing to become a fan and give a medal to whoever can help me with this question

OpenStudy (anonymous):

Colby and Danielle clean pools for extra money over the summer. Colby's income is determined by f(x) = 3x + 12, where x is the number of hours. Danielle's income is g(x) = 5x + 10. If Colby and Danielle were to combine their efforts, their income would be h(x) = f(x) + g(x). Assume Colby works 4 hours. Create the function h(x) and indicate if Colby will make more money working alone or by teaming with Danielle. (1 point) h(x) = 2x + 2, work alone h(x) = 8x + 22, team with Danielle h(x) = 2x + 2, team with Danielle h(x) = 8x + 22, work alone

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (anonymous):

@Luigi0210 @whygeom043 @rayan.q1993 @Awesome781

OpenStudy (anonymous):

Okay I am not that good at math but I am like 99% sure I can help you solve this and please anyone correct me but I am sure this is how it goes. Okay, first you need to find and write a equation of the combined incomes of Colby f(x) and Danielle g(x) and their new equation will be h(x) and they tell you how you should set up the equation like h(x) = f(x) + g(x), so that is exactly what you do and how you should write it like h(x) = 3x + 12 + 5x + 10 and then you would combine the like terms, so what would you get?

OpenStudy (anonymous):

8x+2

OpenStudy (anonymous):

No, you would get h(x) = 8x + 22 because there are no negatives and you don't need to move any numbers from one side to another so you keep thier identity and just combine them like you did for the x's. Okay so now that you got the equation you need to figure our if Colbly will earn more money alone or with danielle on 4 hours so you would plug in 4 for x for the f(x) equation and for the new one and that would look like Colby alone f(x) = 3 (4) + 12 = ? Colby plus danielle h(x) 8(4) + 22 = ? Which equation does Colby make more money and that is your answer

OpenStudy (anonymous):

colby and danielle

OpenStudy (fibonaccichick666):

ok, if you need me just tag me again

OpenStudy (anonymous):

yes, I believe so, hoped I helped some !

OpenStudy (anonymous):

yes you did

OpenStudy (anonymous):

Nice, good luck!:) and remember you are smart!!

OpenStudy (anonymous):

Tooottles